# Diffusion Model in Hyperspectral Image Processing and Analysis: A Review

### v1

@@ VI-A2 Experimental results
Table XVIII is roughly in agreement with Tables XV, XVI and XVII diffusion models method noise suppression performance is generally the best, and in combination with the above tables, the overall effect is highly acceptable, with obvious denoising effect, close to noiseless, and with better detail retention, despite the fact that the temporal processing is not particularly fast.
VI-B Key discoveries and Summaries from the Experiments
@@
Through the experimental comparisons in Section V.B.2, the diffusion model shows significant advantages in hyperspectral image processing tasks, but also exposes some limitations.
@@
In the classification task, the OA (93.15%, 94.78%, 98.97%), AA (96.44%, 93.84%, 99.47%) and Kappa coefficients of the SpectralDiff method based on the diffusion model on the Indian Pines, Pavia University and Salinas datasets ( 92.18%, 93.06%, 98.85%) are better than traditional deep learning methods (e.g., GAN, VAE), indicating that it can effectively capture the complex spectral-spatial features of hyperspectral data. However, its inference time (48.53 s, 256.23 s, 242.16 s) is significantly higher than that of lightweight models such as miniGCN (0.95 s, 2.40 s, 2.13 s), revealing the computational efficiency bottleneck of the diffusion model in high-dimensional data processing.
@@
In the target detection task, the diffusion-based Diff-Mosaic method achieves optimal IoU (79.44%, 91.18%), detection probability (Pd: 99.99%, 99.47%), and false alarm rate (Fa: 3.19%, 1.91%) on the SIRST and NUDT-SIRST datasets, and is especially good at complex background suppression and target localization accuracy. However, the number of parameters (4.7M) and the inference time (43.42 s) are still higher than that of the traditional method ACM (0.29M, 18.53 s), which indicates that the diffusion model needs to be further lightweighted to meet the real-time demand.
@@
In the anomaly detection task, DWSDiff comprehensively leads in mAUC metrics, such as mAUC(D,F) (0.97427), mAUC(BS) (0.95791), and mAUC(SNPR) (67.20618), which verifies its ability to suppress complex backgrounds through the dual-window strategy and iterative diffusion. However, its inference time (0.8094 s) is still much higher than that of Auto-AD (0.0022 s), and the model is sensitive to hyperparameters (e.g., noise scheduling), and its generalization ability needs to be further optimized, suggesting that the practical deployment of diffusion models in anomaly detection needs to balance accuracy and efficiency.
@@
In image recovery tasks, diffusion model methods (e.g., FastDiffSR, EDiffSR, and DDPM) significantly outperform conventional methods in key metrics (PSNR, FID, LPIPS). For example, on the Potsdam dataset, FastDiffSR achieves a PSNR of 33.13 dB and a FID of 22.83, indicating that it outperforms conventional methods in both detail fidelity and distribution consistency. In addition, FastDiffSR achieves a PSNR of 25.53 dB and a FID of 55.76 on the Toronto dataset, which further validates its advantage in semantic information retention. However, the high number of parameters of the diffusion model method (e.g., 23.23M for FastDiffSR) leads to a significant increase in training time. It shows that the diffusion model has strong generative capability, but its computational efficiency and model complexity still need to be optimized to meet the actual deployment requirements.
@@
In noise suppression tasks, diffusion model-based methods (e.g., Diff-Unmix, DDRM, and DDS2M) show significant advantages in metrics such as PSNR, SSIM, and FSIM. For example, on the KAIST dataset, Diff-Unmix achieves a PSNR of 31.41 dB, an SSIM of 0.902, and an FSIM of 0.958, indicating significant advantages in noise suppression, structure preservation, and spectral fidelity. In addition, on the CAVE dataset, Diff-Unmix has a PSNR of 32.71 dB and an SSIM of 0.940, further validating its superiority. However, the long inference time of the diffusion model method (e.g., 42 s for Diff-Unmix) restricts its application in scenarios with high real-time requirements, suggesting that the diffusion model is able to remove the noise and retain the details efficiently, but its computational efficiency still needs to be further improved.
@@ VII Outlook
VIII Conclusion
In summary, compared with other deep generative models, the diffusion model is robust to the inherent noise in hyperspectral images, better adapts to its variability and complexity, and provides a more stable training process. Meanwhile, diffusion model with progressive denoising mechanism with high-dimensional data modeling capability, the diffusion model has injected a revolutionary impetus into the field of hyperspectral image analysis. It effectively solves the limitations of traditional methods in high-dimensional data processing, noise interference and small sample scenarios through the Markov chain noise injection and inverse generation process. Therefore, it is an inevitable trend for diffusion model to be applied in the field of hyperspectral images. This paper firstly introduces the related knowledge of hyperspectral as well as the basic theory of diffusion model, and then, reviews the research of diffusion model in the field of hyperspectral images, including: data enhancement, generation and complementation, classification segmentation, anomaly detection, target detection, denoising and other tasks, moreover, this paper fully demonstrates the superiority of diffusion model in this field by comparing some visualization results and quantitative indexes on hyperspectral image processing between diffusion model and traditional model. The limitations of the existing diffusion models in the hyperspectral domain are also discussed, and it is emphasized that further exploratory studies can be carried out for extended applications and model deployment.

### v2

@@ 6.1.2 Experimental results
Table 20 largely corroborates the findings from Tables 17, 18, and 19, confirming that diffusion model-based methods generally offer the best noise suppression performance. When considering all these tables together, the overall effect of diffusion models is highly acceptable, demonstrating a clear denoising effect that closely approaches noiseless images while retaining superior detail. This is true despite their processing time not being exceptionally fast.
6.2 Key discoveries and Summaries from the Experiments
@@
As evidenced by the experimental comparisons in Section 5.2.2, diffusion models exhibit significant advantages in hyperspectral image processing tasks, while also revealing certain limitations.
@@
In classification tasks, the SpectralDiff method, based on diffusion models, demonstrates superior accuracy metrics (OA, AA, and Kappa coefficients) across the Indian Pines, Pavia University, and Salinas datasets. For instance, its OA values were 93.15%, 94.78%, and 98.97% respectively, outperforming traditional deep learning methods like GANs and VAEs. This indicates its effectiveness in capturing complex spectral-spatial features. However, a significant drawback is its inference time, which is considerably higher (e.g., 48.53 s on Indian Pines) compared to lightweight models such as miniGCN (0.95 s), exposing a computational efficiency bottleneck for high-dimensional data processing.
@@
For target detection, the diffusion-based Diff-Mosaic method achieves optimal performance on the SIRST and NUDT-SIRST datasets, with excellent IoU (79.44% on SIRST), high detection probability (99.99% on SIRST), and a low false alarm rate (3.19% on SIRST). It particularly excels at complex background suppression and precise target localization. Despite this, its parameter count (4.7M) and inference time (43.42 s) remain higher than traditional methods like ACM (0.29M parameters, 18.53 s inference time). This suggests that further lightweighting of diffusion models is needed to meet real-time demands in target detection.
@@
In anomaly detection, DWSDiff comprehensively leads in mAUC metrics (e.g., ${mAUC}_{TD}$ of 0.97427), validating its ability to suppress complex backgrounds through its dual-window strategy and iterative diffusion. While BSDM’s performance is also respectable, DWSDiff stands out. Nevertheless, DWSDiff’s inference time (0.8094 s) is still substantially higher than highly efficient methods like Auto-AD (0.0022 s). Additionally, the model’s sensitivity to hyperparameters (e.g., noise scheduling) and its need for further optimization in generalization ability indicate that practical deployment of diffusion models in anomaly detection requires a careful balance between accuracy and efficiency.
@@
For image recovery, diffusion model methods such as FastDiffSR, EDiffSR, and DDPM significantly outperform conventional methods in key metrics like PSNR and FID. For example, FastDiffSR achieved a PSNR of 33.13 dB and an FID of 22.83 on the Potsdam dataset, demonstrating superior detail fidelity and distributional consistency. Similarly, on the Toronto dataset, FastDiffSR maintained its advantage with a PSNR of 25.53 dB and an FID of 55.76, confirming its strength in semantic information retention. However, the high number of parameters in these diffusion models (e.g., FastDiffSR with 23.23M parameters) leads to a substantial increase in training time. This highlights that while diffusion models possess strong generative capabilities, their computational efficiency and model complexity require further optimization for practical deployment.
@@
In noise suppression tasks, diffusion model-based methods, including Diff-Unmix, DDRM, and DDS2M, show significant advantages across metrics like PSNR, SSIM, and FSIM. On the KAIST dataset, Diff-Unmix achieved a PSNR of 31.41 dB, an SSIM of 0.902, and an FSIM of 0.958, underscoring its benefits in noise suppression, structure preservation, and spectral fidelity. Its superiority was further validated on the CAVE dataset with a PSNR of 32.71 dB and an SSIM of 0.940. Despite their effectiveness in removing noise and retaining details, the long inference time of diffusion models (e.g., 42 s for Diff-Unmix) restricts their application in scenarios demanding high real-time performance, suggesting a need for improved computational efficiency.
@@ 7 Outlook
8 Conclusion
In summary, diffusion models offer significant advantages over other deep generative models for hyperspectral image analysis. They demonstrate robustness to the inherent noise in hyperspectral images, adapt better to their variability and complexity, and provide a more stable training process. With their progressive denoising mechanism and high-dimensional data modeling capabilities, diffusion models have injected a revolutionary impetus into the field of hyperspectral image analysis. They effectively overcome the limitations of traditional methods in handling high-dimensional data, noise interference, and small sample scenarios through their Markov chain noise injection and inverse generation process. Therefore, the application of diffusion models in hyperspectral imaging is an inevitable trend. This paper first introduced relevant knowledge about hyperspectral images and the basic theory of diffusion models. Subsequently, it reviewed research on diffusion models within the hyperspectral domain, covering tasks such as data enhancement, generation and complementation, classification segmentation, anomaly detection, target detection, and denoising. Moreover, by comparing various visualization results and quantitative indices between diffusion models and traditional models in hyperspectral image processing, this paper fully demonstrated the superiority of diffusion models in this field. Finally, the limitations of existing diffusion models in the hyperspectral domain were discussed, emphasizing the need for further exploratory studies concerning extended applications and model deployment.